# Anisotropic Green Coordinates

### v1

@@ 4. Derivation of Closed-form Anisotropic Green Coordinates
In previous sections, we derived the deformation formula for anisotropic Green coordinates and obtained the integral expressions for $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ in Eqs. (24) and (25), respectively. Additionally, we explored several properties associated with this deformation. In this section, we derive the closed-form expressions for $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ in both 2D and 3D scenarios. These analytical solutions are more accurate and computationally efficient compared to numerical methods for evaluating the integrals. In the following derivations, we frequently use formulations for the substitution of area elements on the level-set manifold within embedded simplices. These formulations are provided as Lemmas A.3, A.4, and A.5 in the supplementary materials. Although these lemmas do not rely on advanced mathematical concepts, they are not entirely trivial or immediately obvious. For this reason, detailed proofs are also included in the supplementary materials. Moreover, in Section B of the supplementary material, we provide a detailed derivation of the closed-form expressions for isotropic Green coordinates, including their gradients and Hessians, which form the basis for the anisotropic formulation.
@@ 4.1. Closed-form Expressions for 2D Scenario
With the above lemmas, we can now derive the closed-form expressions for the Dirichlet term $\phi_{i}^{\mathbf{A}}(\eta)$ and Neumann term $\psi_{j}^{\mathbf{A}}(\eta)$ of the anisotropic Green coordinates. We begin by presenting the method in 2D. We assume that the 2D cage is a closed and oriented 1-mainfold in $\mathbb{R}^{2}$ with vertices $I_{\mathbb{V}}=\{\mathbf{v_{i}}\}_{i=1}^{N}$ and faces $I_{\mathbb{T}}=\{f_{j}|f_{j}=\overrightarrow{\mathbf{v}_{j}\mathbf{v}_{j+1}}\}$ arranged in counter-clockwise order. We denote $\mathbf{a}_{j}=\mathbf{v}_{j+1}-\mathbf{v}_{j}$. Then, the outer normal vector $\mathbf{n}_{j}$ of $f_{j}$ is then given by $\mathbf{n}_{j}={\mathbf{a}_{j}}^{\bot}/\|{\mathbf{a}_{j}}^{\bot}\|$, where the superscript $\bot$” denotes the vector obtained by rotating the original vector 90 degrees clockwise. Then, $\psi_{j}^{\mathbf{A}}(\eta)$ has the following expression:
@@
The key step is to transform the anisotropic calculation into the isotropic one through a variable substitution. Let $\mathbf{x}=\mathbf{A}^{-1/2}\xi,\ \mathbf{y}=\mathbf{A}^{-1/2}\eta$, $\mathbf{v}^{\prime}_{j}=\mathbf{A}^{-1/2}\mathbf{v}_{j}$ and $\mathbf{v}^{\prime}_{j+1}=\mathbf{A}^{-1/2}\mathbf{v}_{j+1}$. According to Lemma A.3 and Lemma A.4 of the supplementary material, we have $\mathrm{d}\sigma_{\mathbf{x}}=|\det\mathbf{A}^{-1/2}|\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|\ \mathrm{d}\sigma_{\xi}$, where $\mathbf{n}_{j}$ is the unit normal of $f_{j}$. Denote $f_{j}^{\prime}=\overrightarrow{\mathbf{v}_{j}^{\prime}\mathbf{v}_{j+1}^{\prime}}$, we have
@@
(33) $=$ $\frac{1}{|\det{\mathbf{A}^{-1/2}}|\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|}\int_{\xi\in f_{j}^{\prime}}-{\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}\log\|\mathbf{x}-\mathbf{y}\|}\ \mathrm{d}\sigma_{\mathbf{x}}$
(34) $=$ $\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\int_{\xi\in f_{j}^{\prime}}{-\frac{1}{2\pi}\log\|\mathbf{x}-\mathbf{y}\|}\ \mathrm{d}\sigma_{\mathbf{x}}.$
@@
Next, we describe the calculation of the Dirichlet term $\phi_{\mathbf{v}_{j}}^{\mathbf{A}}(\eta)$ for the anisotropic Green coordinates. Since $\phi_{\mathbf{v}_{j}}^{\mathbf{A}}(\eta)$ is obtained by summing the integrals over the edges associated with $\mathbf{v}_{i}$, we focus on the integral over a single edge $f_{j}$, which is denoted as $\phi_{\mathbf{v}_{j},f_{j}}^{\mathbf{A}}(\eta)$. Recall that $\mathbf{x}=\mathbf{A}^{-1/2}\xi,\ \mathbf{y}=\mathbf{A}^{-1/2}\eta$, $\mathbf{v}^{\prime}_{j}=\mathbf{A}^{-1/2}\mathbf{v}_{j}$, $\mathbf{v}^{\prime}_{j+1}=\mathbf{A}^{-1/2}\mathbf{v}_{j+1}$, $f_{j}=\overrightarrow{\mathbf{v}_{j}\mathbf{v}_{j+1}}$ and $f_{j}^{\prime}=\overrightarrow{\mathbf{v}_{j}^{\prime}\mathbf{v}_{j+1}^{\prime}}$. Let $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi)$ be the hat-function of $\xi$ toward $\mathbf{v}_{j}$ in $f_{j}$, and $\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x})$ be the hat-function of $\mathbf{x}$ toward $\mathbf{v}_{j}^{\prime}$ in $f_{j}^{\prime}$. We first show that $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})=\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0})$ for $\xi_{0}\in f_{j}$ and $\mathbf{x}_{0}=\mathbf{A}^{-1/2}\xi_{0}$. First, we have
@@
(37) $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi)+\Gamma_{\mathbf{v}_{j+1},f_{j}}(\xi)=1.$
Therefore,
@@
Therefore, $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0}),\Gamma_{\mathbf{v}_{j+1},f_{j}}(\xi_{0})$ and $\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0}),\Gamma_{\mathbf{v}_{j+1}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0})$ are both barycentric coordinates of $\mathbf{x}_{0}$ in $f_{j}^{\prime}$. As established in Lemma A.6 of the supplementary material, the barycentric coordinates of a simplex are uniquely determined due to the affine independence of its vertices. This property distinguishes them from barycentric coordinates in arbitrary polygonal domains, where uniqueness is not guaranteed. Therefore, we conclude that $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})=\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0})$ for any $\xi_{0}\in f_{j}$ and $\mathbf{x}_{0}=\mathbf{A}^{-1/2}\xi_{0}$. Moreover, we have $\mathrm{d}\sigma_{\xi}=|\det\mathbf{A}^{1/2}|\|\mathbf{A}^{-1/2}\mathbf{n}_{j}^{\prime}\|\mathrm{d}\sigma_{\mathbf{x}}$, where $\mathbf{n}_{j}^{\prime}$ is the unit normal of $\overrightarrow{\mathbf{v}_{j}^{\prime}\mathbf{v}_{j+1}^{\prime}}$. Then, we can derive the closed-from solution of $\phi_{\mathbf{v}_{j},f_{j}}^{\mathbf{A}}(\eta)$ as follows:
@@
Recall that $\mathbf{a}_{j}=\mathbf{v}_{j+1}^{\prime}-\mathbf{v}_{j}^{\prime}$ and $\mathbf{b}_{j}=\mathbf{v}_{j}^{\prime}-\mathbf{y}$. The term
@@ 4.2. Closed-form Expressions for 3D Scenario
We can employ a similar methodology to derive the closed-form expression for anisotropic Green coordinates in three dimensions. Fig. 3 illustrates a tetrahedron formed by a triangular face $t_{j}=\triangle\mathbf{v}_{j_{1}}\mathbf{v}_{j_{2}}\mathbf{v}_{j_{3}}$ of the cage and an interior point $\eta$ located within the cage. Let $\mathbf{x}=\mathbf{A}^{-1/2}\xi,\ \mathbf{y}=\mathbf{A}^{-1/2}\eta$, $\mathbf{v}^{\prime}_{j_{1}}=\mathbf{A}^{-1/2}\mathbf{v}_{j_{1}}$, $\mathbf{v}^{\prime}_{j_{2}}=\mathbf{A}^{-1/2}\mathbf{v}_{j_{2}}$ and $\mathbf{v}^{\prime}_{j_{3}}=\mathbf{A}^{-1/2}\mathbf{v}_{j_{3}}$. Based on Lemma A.5 in the supplementary material, and given that $\mathbf{A}$ is symmetric, we have $\mathrm{d}\sigma_{\mathbf{x}}=|\det\mathbf{A}^{-1/2}|\|\mathbf{A}^{1/2}\mathbf{n}_{t}\|\ \mathrm{d}\sigma_{\xi}$, where $\mathbf{n}_{t}$ represents the unit normal vector of $t_{j}$. Moreover, denoting $t_{j}^{\prime}=\triangle\mathbf{v}_{j_{1}}^{\prime}\mathbf{v}_{j_{2}}^{\prime}\mathbf{v}_{j_{3}}^{\prime}$ with normal $\mathbf{n}_{t}^{\prime}$. Then, the Neumann term $\psi_{j}^{\mathbf{A}}(\eta)$ of $t_{j}$ can be computed as:
@@
For the Dirichlet term, we apply $\phi_{\mathbf{v}_{j_{1}},t_{j}}^{\mathbf{A}}(\eta)$ to denote the contribution of $t_{j}$ to $\phi_{\mathbf{v}_{j_{1}}}^{\mathbf{A}}(\eta)$. According to the uniqueness of the barycentric coordinates of a simplex (Lemma A.6), we also have $\Gamma_{\mathbf{v}_{j_{1}},t_{j}}(\xi_{0})=\Gamma_{\mathbf{v}_{j_{1}}^{\prime},t_{j}^{\prime}}(\mathbf{x}_{0})$ for $\xi_{0}\in t_{j}$ and $\mathbf{x}_{0}=\mathbf{A}^{-1/2}\xi_{0}$. Therefore,
@@
Then, we convert the computation of anisotropic coordinates into isotropic coordinates. The calculation of 3D isotropic Green coordinates is relatively complex, with details provided in Section B.2 of the supplementary material. The approach is primarily derived from the findings in (2009VariationalGC) and (Urago2000), but we offer a more detailed derivation.
@@ 5. Discussion
In the previous section, we derived the expression for anisotropic Green coordinates and obtained their closed-form representation. In this section, we will clarify the geometric interpretation of anisotropic Green coordinates in comparison to the isotropic counterparts and present a proof. Furthermore, we will emphasize the essential value of considering problems from an anisotropic perspective. We have the following observation when ignoring the effect of the scale factor $s_{j}$: the two approaches below are equivalent: 1. Given an input, first multiply the input globally, including the source cage, target cage, and the object, by $\mathbf{A}^{-1/2}$, then apply the isotropic deformation, and finally multiply globally by $\mathbf{A}^{1/2}$; 2. Using the anisotropic Green coordinates for the source and target cage. However, we believe that there are still important points to illustrate the value of deriving the anisotropic Green coordinates. First, our method provides a geometric interpretation of the integral formula in terms of the anisotropic Laplacian equation. Second, as deformation coordinates, they need to be expressed as functions relative to the original cage, rather than the cage has undergone a global transformation. Finally, when computing the Jacobian and Hessian of the coordinates, and applying the variational method for deformation control—which will be presented in Section 6—we require the expressions of $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{i}^{\mathbf{A}}(\eta)$ in terms of the anisotropic Green coordinates.
Next, we refer to the deformation effect obtained by “globally multiplying by $\mathbf{A}^{-1/2}$, performing deformation using isotropic Green coordinates, and then globally multiplying by $\mathbf{A}^{1/2}$”, as similarity Green deformation. The name stems from the fact that this process closely resembles a matrix similarity transformation. Subsequently, we prove that anisotropic Green deformation is equivalent to similarity Green deformation. We denote the vertices and normals of the source and target cage as $\{\mathbf{v}_{i}\}_{i=1}^{N}$ and $\{\mathbf{n}_{j}\}_{j=1}^{M}$. Moreover, we use the prime operator $\prime$” to denote the globally transformed values, e.g., $\mathbf{v}_{i}^{\prime}=\mathbf{A}^{-1/2}\mathbf{v}_{i}$, $\mathbf{n}_{j}^{\prime}=\frac{\mathbf{A}^{1/2}\mathbf{n}_{j}}{\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|}$. Then, we can express the globally transformed values in terms of isotropic Green coordinates as:
@@
The derivation of section 4.2 shows that $\phi_{i}^{\prime}(\eta^{\prime})=\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\prime}(\eta^{\prime})=\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|\psi_{j}^{\mathbf{A}}(\eta)$, where $\mathbf{n}_{j}$ is the normal of $f_{j}$. Moreover, we have $\eta=\mathbf{A}^{1/2}\eta^{\prime}$. Then,
@@
and we complete the proof of the equivalence between the anisotropic Green deformation and the similarity Green deformation.

### v2

@@ 4. Derivation of Closed-form Anisotropic Green Coordinates
We have derived the deformation formula for anisotropic Green coordinates and investigated its fundamental properties. In this section, we present closed-form expressions for $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ in both 2D and 3D. These analytical solutions are more accurate and computationally efficient compared to numerical methods for evaluating the integrals. In the following derivations, we frequently use formulations for the substitution of area elements on curves and surfaces embedded in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. These formulations are detailed in Lemmas A.4, A.5, and A.6 of the supplementary material. While these results do not rely on advanced mathematical concepts, they are non-trivial and not immediately obvious. Additionally, Appendix B provides a detailed derivation of the closed-form expressions for classical isotropic Green coordinates and their gradients and Hessians, which serve as the foundation for the anisotropic formulation.
@@ 4.1. Closed-form Expressions for 2D Scenario
We begin by presenting the method in 2D. We assume that the 2D cage is a closed and oriented 1-mainfold in $\mathbb{R}^{2}$ with vertices $I_{\mathbb{V}}=\{\mathbf{v_{i}}\}_{i=1}^{N}$ and faces $I_{\mathbb{T}}=\{f_{j}|f_{j}=\overrightarrow{\mathbf{v}_{j}\mathbf{v}_{j+1}}\}$ arranged in counter-clockwise order. We denote $\mathbf{a}_{j}=\mathbf{v}_{j+1}-\mathbf{v}_{j}$. Then, the unit outward normal $\mathbf{n}_{j}$ of $f_{j}$ is given by $\mathbf{n}_{j}={\mathbf{a}_{j}}^{\bot}/\|{\mathbf{a}_{j}}^{\bot}\|$, where the superscript $\bot$” denotes the vector obtained by rotating the original vector 90 degrees clockwise. Then, $\psi_{j}^{\mathbf{A}}(\eta)$ has the following expression:
@@
The key step is to transform the calculation of the anisotropic formulation into the isotropic one through the variable substitution. Specifically, let $\mathbf{x}=\mathbf{A}^{-1/2}\xi,\ \mathbf{y}=\mathbf{A}^{-1/2}\eta$, $\mathbf{v}^{\prime}_{j}=\mathbf{A}^{-1/2}\mathbf{v}_{j}$ and $\mathbf{v}^{\prime}_{j+1}=\mathbf{A}^{-1/2}\mathbf{v}_{j+1}$. According to Lemma A.4 and Lemma A.5 of the supplementary material, we have $\mathrm{d}\sigma_{\mathbf{x}}=|\det\mathbf{A}^{-1/2}|\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|\ \mathrm{d}\sigma_{\xi}$, where $\mathbf{n}_{j}$ is the unit normal of $f_{j}$. Denoting $f_{j}^{\prime}=\overrightarrow{\mathbf{v}_{j}^{\prime}\mathbf{v}_{j+1}^{\prime}}$, we have
@@
$=$ $\frac{1}{|\det{\mathbf{A}^{-1/2}}|\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|}\int_{\xi\in f_{j}^{\prime}}-{\frac{1}{2\pi\sqrt{\det{\mathbf{A}}}}\log\|\mathbf{x}-\mathbf{y}\|}\ \mathrm{d}\sigma_{\mathbf{x}}$
$=$ $\frac{1}{\sqrt{\mathbf{n}_{j}^{\top}\mathbf{A}\mathbf{n}_{j}}}\int_{\xi\in f_{j}^{\prime}}{-\frac{1}{2\pi}\log\|\mathbf{x}-\mathbf{y}\|}\ \mathrm{d}\sigma_{\mathbf{x}}.$
@@
Please refer to Appendix B.1 for more details on this formula.
We now proceed to the Dirichlet term $\phi_{\mathbf{v}_{j}}^{\mathbf{A}}(\eta)$, which is obtained by summing the contributions over the edges associated with $\mathbf{v}_{i}$. We focus on the integral over a single edge $f_{j}$, denoted as $\phi_{\mathbf{v}_{j},f_{j}}^{\mathbf{A}}(\eta)$. Recall that $\mathbf{x}=\mathbf{A}^{-1/2}\xi,\ \mathbf{y}=\mathbf{A}^{-1/2}\eta$, $\mathbf{v}^{\prime}_{j}=\mathbf{A}^{-1/2}\mathbf{v}_{j}$, $\mathbf{v}^{\prime}_{j+1}=\mathbf{A}^{-1/2}\mathbf{v}_{j+1}$, $f_{j}=\overrightarrow{\mathbf{v}_{j}\mathbf{v}_{j+1}}$ and $f_{j}^{\prime}=\overrightarrow{\mathbf{v}_{j}^{\prime}\mathbf{v}_{j+1}^{\prime}}$. Let $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi)$ be the hat-function of $\xi$ with respect to $\mathbf{v}_{j}$ in $f_{j}$, and $\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x})$ be the hat-function of $\mathbf{x}$ with respect to $\mathbf{v}_{j}^{\prime}$ in $f_{j}^{\prime}$. We first show that $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})=\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0})$ for $\xi_{0}\in f_{j}$ and $\mathbf{x}_{0}=\mathbf{A}^{-1/2}\xi_{0}$. Recall that $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})$ is the barycentric coordinates of $\xi_{0}$ with respect to $\mathbf{v}_{j}$ in $f_{j}$, we obtain
@@
(35) $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})+\Gamma_{\mathbf{v}_{j+1},f_{j}}(\xi_{0})=1.$
Then,
@@
Therefore, $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0}),\Gamma_{\mathbf{v}_{j+1},f_{j}}(\xi_{0})$ and $\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0}),\Gamma_{\mathbf{v}_{j+1}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0})$ are both barycentric coordinates of $\mathbf{x}_{0}$ in $f_{j}^{\prime}$. As established in Lemma A.7, the barycentric coordinates of a simplex are uniquely determined due to the affine independence of its vertices. This property distinguishes them from barycentric coordinates in arbitrary polygonal domains, where uniqueness is generally not guaranteed. Therefore, we conclude that $\Gamma_{\mathbf{v}_{j},f_{j}}(\xi_{0})=\Gamma_{\mathbf{v}_{j}^{\prime},f_{j}^{\prime}}(\mathbf{x}_{0})$ for any $\xi_{0}\in f_{j}$ and $\mathbf{x}_{0}=\mathbf{A}^{-1/2}\xi_{0}$. Moreover, we have $\mathrm{d}\sigma_{\xi}=|\det\mathbf{A}^{1/2}|\|\mathbf{A}^{-1/2}\mathbf{n}_{j}^{\prime}\|\mathrm{d}\sigma_{\mathbf{x}}$, where $\mathbf{n}_{j}^{\prime}$ is the unit normal of $\overrightarrow{\mathbf{v}_{j}^{\prime}\mathbf{v}_{j+1}^{\prime}}$. Then, we can derive the closed-from solution of $\phi_{\mathbf{v}_{j},f_{j}}^{\mathbf{A}}(\eta)$ as follows:
@@
Denote $\mathbf{a}_{j}=\mathbf{v}_{j+1}^{\prime}-\mathbf{v}_{j}^{\prime}$ and $\mathbf{b}_{j}=\mathbf{v}_{j}^{\prime}-\mathbf{y}$. The term
@@ 4.2. Closed-form Expressions for 3D Scenario
We can employ a similar methodology to derive the closed-form expression for the 3D case. Figure 3 depicts a tetrahedron formed by a triangular face $t_{j}=\triangle\mathbf{v}_{j_{1}}\mathbf{v}_{j_{2}}\mathbf{v}_{j_{3}}$ of the cage and an interior point $\eta$ within the cage. Let $\mathbf{x}=\mathbf{A}^{-1/2}\xi,\ \mathbf{y}=\mathbf{A}^{-1/2}\eta$, $\mathbf{v}^{\prime}_{j_{1}}=\mathbf{A}^{-1/2}\mathbf{v}_{j_{1}}$, $\mathbf{v}^{\prime}_{j_{2}}=\mathbf{A}^{-1/2}\mathbf{v}_{j_{2}}$ and $\mathbf{v}^{\prime}_{j_{3}}=\mathbf{A}^{-1/2}\mathbf{v}_{j_{3}}$. According to Lemma A.6, we have $\mathrm{d}\sigma_{\mathbf{x}}=|\det\mathbf{A}^{-1/2}|\|\mathbf{A}^{1/2}\mathbf{n}_{t}\|\ \mathrm{d}\sigma_{\xi}$, where $\mathbf{n}_{t}$ denotes the unit outward normal vector of $t_{j}$. Furthermore, denoting $t_{j}^{\prime}=\triangle\mathbf{v}_{j_{1}}^{\prime}\mathbf{v}_{j_{2}}^{\prime}\mathbf{v}_{j_{3}}^{\prime}$, and its normal as $\mathbf{n}_{t}^{\prime}$. Then, the Neumann term $\psi_{j}^{\mathbf{A}}(\eta)$ on $t_{j}$ can be computed as follows:
@@
For the Dirichlet term, we denote by $\phi_{\mathbf{v}_{j_{1}},t_{j}}^{\mathbf{A}}(\eta)$ the contribution of $t_{j}$ to $\phi_{\mathbf{v}_{j_{1}}}^{\mathbf{A}}(\eta)$. According to Lemma A.6, we have $\mathrm{d}\sigma_{\xi}=|\det\mathbf{A}^{1/2}|\|\mathbf{A}^{-1/2}\mathbf{n}_{t}^{\prime}\|\mathrm{d}\sigma_{\mathbf{x}}$. Based on the uniqueness of the barycentric coordinates of a simplex (Lemma A.7), it holds that $\Gamma_{\mathbf{v}_{j_{1}},t_{j}}(\xi_{0})=\Gamma_{\mathbf{v}_{j_{1}}^{\prime},t_{j}^{\prime}}(\mathbf{x}_{0})$ for $\xi_{0}\in t_{j}$ and $\mathbf{x}_{0}=\mathbf{A}^{-1/2}\xi_{0}$. Therefore,
@@
Then, we transform the computation of the anisotropic formulation into an isotropic one. The calculation of 3D isotropic Green coordinates is somewhat complex, and we provide the details in Appendix B.2 of the supplementary material. These derivations are primarily based on the findings of  (Ben-Chen et al., 2009) and  (Urago, 2000), but we provide a more comprehensive derivation.
@@ 5. Discussion
In this section, we mainly derive the geometric interpretation of anisotropic Green coordinates in comparison to the isotropic counterparts, and demonstrate the quasi-conformality of our method. Furthermore, we emphasize the essential value of addressing problems from an anisotropic perspective.
We have the following observation when ignoring the effect of the scale factor $s_{j}$: the two approaches below are equivalent:
  1. (1)
    Given an input, first apply a global multiplication by $\mathbf{A}^{-1/2}$ to the source cage, target cage, and the object; subsequently deform the object using the isotropic Green coordinates; and finally multiply globally by $\mathbf{A}^{1/2}$.
  2. (2)
    Employ the anisotropic Green coordinates for the source and target cages.
We will present the derivation of this equivalence subsequently. Despite the equivalence, we believe that there are still important points to illustrate the value of deriving the anisotropic Green coordinates. First, our method provides a geometric interpretation of Green’s identity derived from the anisotropic Laplace equation. Second, as deformation coordinates, they need to be expressed as functions relative to the original cage, rather than the cage has undergone a global transformation. Moreover, when performing the variational method for deformation control (which will be presented in Section 6), it is more intuitive to analyze directly from the closed-form anisotropic formulation as well as its gradient and Hessian. Most importantly, during user interaction, the deformation coordinates $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ typically need to be calculated only once for the input source cage, and subsequent deformations do not require recalculation. If calculated directly using anisotropic Green coordinates, according to Eq. (26), only $s_{j}$, $\mathbf{A}\tilde{\mathbf{n}}_{j}$ and $\tilde{\eta}$ are required to be computed every time. In contrast, if calculated from the perspective of the global transform, it is necessary to first compute the transformed target cage and normals, which increases the computational overhead. Furthermore, this study introduces a new perspective on cage-based and variational deformation. If the matrix $\mathbf{A}$ can be optimized, there is potential to further enhance the practical value of the method, which remains a direction for our future work.
Next, we refer to the deformation effect obtained by “globally multiplying by $\mathbf{A}^{-1/2}$, performing deformation using isotropic Green coordinates, and then globally multiplying by $\mathbf{A}^{1/2}$”, as similarity Green deformation. The name stems from the fact that this process closely resembles a matrix similarity transformation. We denote the vertices and normals of the cage as $\{\mathbf{v}_{i}\}_{i=1}^{N}$ and $\{\mathbf{n}_{j}\}_{j=1}^{M}$. Moreover, we use the prime operator to denote the globally transformed values, e.g., $\mathbf{v}_{i}^{\prime}=\mathbf{A}^{-1/2}\mathbf{v}_{i}$, $\mathbf{n}_{j}^{\prime}=\frac{\mathbf{A}^{1/2}\mathbf{n}_{j}}{\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|}$. Then, we can express the globally transformed values in terms of isotropic Green coordinates as:
@@
The derivation of section 4.2 (Eqs. (32) and (37)) shows that $\phi_{i}^{\prime}(\eta^{\prime})=\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\prime}(\eta^{\prime})=\|\mathbf{A}^{1/2}\mathbf{n}_{j}\|\psi_{j}^{\mathbf{A}}(\eta)$, where $\mathbf{n}_{j}$ is the normal of $f_{j}$. Moreover, we have $\eta=\mathbf{A}^{1/2}\eta^{\prime}$. Then,
@@
and we show the equivalence between the anisotropic Green deformation and the similarity Green deformation.
Moreover, we can use the above analysis to demonstrate the quasi-conformality of our method. We denote the anisotropic deformation mapping as $f_{\mathbf{A}}(\eta)$, and the corresponding isotropic mapping as $f^{\prime}(\eta)$. Then,
(44) $f_{\mathbf{A}}(\eta)=\mathbf{A}^{1/2}f^{\prime}(\mathbf{A}^{-1/2}\eta).$
The Jacobian of $f_{\mathbf{A}}(\eta)$ is given by:
(45) $\mathbf{J}_{f_{\mathbf{A}}}(\eta)=\mathbf{A}^{1/2}\mathbf{J}_{f^{\prime}}(\mathbf{A}^{-1/2}\eta)\mathbf{A}^{-1/2},$
where $\mathbf{J}_{f^{\prime}}$ is the Jacobian of the isotropic formulation. Let $K_{\mathbf{A}}(\eta)$ denote the ratio of the maximum to the minimum singular values of the anisotropic deformation mapping, and $K^{\prime}(\eta)$ the corresponding isotropic ratio. Then,
(46) $K_{\mathbf{A}}(\eta)=$ $\frac{\sigma_{max}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))}{\sigma_{min}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))}=\text{cond}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))$
$=\text{cond}(\mathbf{A}^{1/2}\mathbf{J}_{f^{\prime}}(\mathbf{A}^{-1/2}\eta)\mathbf{A}^{-1/2})$
$\leq\text{cond}(\mathbf{A}^{1/2})\text{cond}(\mathbf{J}_{f^{\prime}}(\mathbf{A}^{-1/2}\eta))\text{cond}(\mathbf{A}^{-1/2})$
$=\text{cond}(\mathbf{A})K^{\prime}(\mathbf{A}^{-1/2}\eta),$
where $\sigma_{max}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))$ and $\sigma_{min}(\mathbf{J}_{f_{\mathbf{A}}}(\eta))$ represent the maximum and minimum singular values of $\mathbf{J}_{f_{\mathbf{A}}}(\eta)$, and $\text{cond}(\mathbf{A})$ represents the condition number of $\mathbf{A}$. Therefore, when Green coordinates are conformal in 2D, anisotropic Green coordinates are quasi-conformal with $K_{\mathbf{A}}(\eta)\leq\text{cond}(\mathbf{A})$. When Green coordinates are quasi-conformal in 3D with a distortion bound $M$, anisotropic Green coordinates are quasi-conformal in 3D with a bound $\text{cond}(\mathbf{A})\cdot M$. Moreover, since $\mathbf{A}=\mathbf{P}^{\top}\Lambda\mathbf{P}$, the condition number $\text{cond}(\mathbf{A})$ is precisely the ratio of the largest to the smallest eigenvalues of $\Lambda$.